OpenStudy (arbyj):

how do i solve 2x^2+11=29 ?

OpenStudy (anonymous):

first take 29 to the other side of the equation

OpenStudy (arbyj):

done

OpenStudy (dayakar):

subtract 11 both sides

OpenStudy (arbyj):

set it up equal to 0 right?

OpenStudy (dayakar):

then divide with 2 both sides

OpenStudy (anonymous):

so now you have 2x^2+11-29=0 2x^2-18=0 2(x^2-9)=0 Factorize it...

OpenStudy (dayakar):

whom ur following arbyj

OpenStudy (arbyj):

lovelarap

OpenStudy (arbyj):

i got 2(x+3)(x-3)

OpenStudy (anonymous):

Your Right! solve for x now

OpenStudy (zenmo):

There is another method of not using "factoring."

OpenStudy (arbyj):

i got -3,3

OpenStudy (zenmo):

You can give the credit to Lovelarap, I'll show you afterwards if you want.

OpenStudy (arbyj):

or x=-3 & x=3

OpenStudy (zenmo):

the order doesn't matter @arbyj

OpenStudy (zenmo):

as long it's x=-3 and x=3.

OpenStudy (arbyj):

@Zenmo factoring is the unit we're learning

OpenStudy (zenmo):

Oh, I see didn't see it in your original post. :)

OpenStudy (anonymous):

yep

OpenStudy (arbyj):

i didn't specify, i apologize

OpenStudy (zenmo):

\[2x^2+11=29\] \[2x^2=18\] \[x^2=9\] \[x = (+/-) \sqrt{9}\] Squaring it will give you two answers: x = -3 & x = 3.

OpenStudy (zenmo):

Another short method

OpenStudy (arbyj):

thank you! @Lovelarap

OpenStudy (anonymous):

your welcome

OpenStudy (arbyj):

thank you @Zenmo (: